Write an equation for a parabola that passes through the given points.
(6, 27), (5, 18), (-4, 27), (-3, 18)

Answer by Boreal(15235)

ax^2+bx+c=y
36a+6b+c=27
25a+5b+c=18
16a-4b+c=27
9a-3b+c=18
-----------------
36a+6b+c=16a-4b+c, because both of these were equal to 27
20a=-10b
b=-2a
------------------
36a-12a+c=27
25a-10a+c=18
24a+c=27
15a+c=18
c=27-24a=18-15a
9a=9
a=1
b=-2
c (from first equation)36-12+c=27, or c=3
Check {1,-2,3} in the other equations ANSWER
25-10+3=18
16+8+3=27
9+6+3=18